# Can Slow-Thinking LLMs Reason Over Time? Empirical Studies in Time Series Forecasting

### v3

@@ 5.3. Implementation Details
The TimeReasoner framework is model-agnostic; in this work, we adopt DeepSeek-R1 for its open-source availability and ability to expose complete reasoning traces. In default settings, all experiments are conducted under the One-Shot Reasoning paradigm, with inference performed via the official API using temperature = 0.6, top-p = 0.7, and max-tokens = 8,192. Foundation model baselines follow the same zero-shot inference protocol, while deep learning baselines are trained using the ADAM optimizer (kingma2014adam) in PyTorch (paszke2019pytorch) on a single NVIDIA GeForce RTX 4090D GPU, with official implementations and recommended parameters. The lookback window is set to $T=36$ for NASDAQ, $T=300$ for VitalDB, and $T=96$ for other datasets; the prediction horizon is $H=36$, $H=200$, and $H=96$ respectively. We adopt a channel-independent evaluation. Following standard TSF practice, we report Mean Squared Error (MSE) to measure squared deviations and Mean Absolute Error (MAE) to measure absolute deviations between predictions and ground truth.
@@ 6.1. Forecasting Performance Comparison
To address RQ1, we evaluate TimeReasoner’s forecasting performance across multiple benchmarks against state-of-the-art baselines. Table 2 summarizes the forecasting results of TimeReasoner, demonstrating its outstanding performance over other baseline models. Results on ten datasets show that TimeReasoner always achieves significant forecasting results, especially on ETTh1 and AQShunyi. The surprising finding is that TimeReasoner particularly excels on datasets characterized by complex temporal dynamics. On datasets like AQWan, AQShunyi and Wind, TimeReasoner consistently surpasses other methods by effectively capturing underlying dependencies. These findings underscore TimeReasoner’s robust capability in tackling difficult forecasting tasks and its generalization to complex time series. It cannot be ignored that performance remains competitive with baselines on simpler datasets. This suggests that while the model excels in modeling complex temporal dynamics, its performance on less intricate time series remains robust yet aligns with established benchmarks.
@@ 6.2. Analysis of Key Settings in TimeReasoner
To address RQ2, we conduct ablations on key configurations, including hybrid prompt components and reasoning strategies, to assess their impact on accuracy.
@@ 6.2.1. Impact of Hybrid Prompt
We assess how different key settings in prompt construction influence TimeReasoner’s forecasting. Accordingly, we investigate how key prompt construction settings—timestamps, contextual information, and normalization—affect TimeReasoner’s forecasting performance (Table 3). For timestamps, we test two perturbations: removal, which discards all timestamp annotations, and shifting, which applies a uniform forward or backward offset while preserving relative intervals. Shifts cause only minor degradation, indicating robustness to absolute-time changes, whereas removal leads to substantial performance drops, confirming the necessity of both absolute and relative temporal cues. For contextual information, removing domain-specific descriptors in prompts yields clear gains on datasets with strong seasonal or location patterns (e.g., Wind) but offers limited benefit on highly volatile series (e.g., NASDAQ), showing that context’s value is domain dependent. For normalization, we compare Z-score and RevIN against using raw data, and find that unnormalized inputs perform best. This suggests that statistical normalization, while often beneficial for deep models, may obscure critical temporal patterns such as trends and abrupt changes, and that preserving the original data distribution better supports slow-thinking reasoning over rich temporal semantics.
Figure 3. Comparison of TimeReasoner’s forecasting capabilities under the three reasoning strategies with experiments conducted on ETTh2 (left) and Exchange (right) datasets.:
@@ 6.2.2. Impact of Various Reasoning Strategies
One hypothesis is that Rollout Reasoning should be more powerful in short-term forecasting because its multi-round reasoning mechanism will have lower cumulative errors than long-term forecasting. In terms of long-term forecasting, Decoupled Reasoning should have more advantages because it emphasizes the reasoning optimization of LLMs in one reasoning. We fix the lookback window $L=96$ and compare the forecasting performance under four predict window lengths $H=\left\{48,96,144,192\right\}$. As shown in Figure 3, One-Shot Reasoning performs well across all predict windows. Rollout Reasoning achieves the best performance at shorter windows (48 and 96). However, as the prediction window extends to 144 and 192, its error increases more rapidly, and Decoupled Reasoning surpasses it to become the best-performing method. This suggests that while One-Shot may struggle with long-range dependencies and Rollout suffers from error accumulation, Decoupled Reasoning benefits from internal refinement, maintaining stability and higher accuracy over longer sequences.
Table 4. Experimental results on ETTh1 and ETTm1 datasets showing model robustness under different input settings.
@@ 6.3. Trustworthiness Evaluation
To address RQ3, we simulate missing and noisy inputs to assess TimeReasoner’s robustness and uncertainty estimation capability.
6.3.1. Robustness Evaluation w.r.t. Missing Input
We evaluate the robustness of TimeReasoner to missing values, a common challenge in real-world time series (e.g., due to sensor failures (che2018recurrent)). While traditional models typically rely on imputation, slow-thinking LLM-based approaches can process raw, incomplete data directly. We consider three variants: (1) No Imputation (No-Imp), where missing entries are removed; (2) None Placeholder Imputation (None-Imp), where missing values are replaced with a None token; and (3) Linear Interpolation Imputation (Lin-Imp) (chow1971best). As shown in Table 4, TimeReasoner performs best with fully observed data, demonstrating its ability to capture complete temporal dependencies. It remains robust under linear interpolation, while using None placeholders causes moderate degradation due to increased uncertainty. In contrast, dropping missing entries leads to a sharp decline, as it disrupts the temporal structure. These results suggest that preserving temporal continuity—even when using placeholders—is more beneficial than deleting missing entries.
Figure 4. Relationship between TimeReasoner’s forecasting performance and lookback window and predict window.:
Figure 5. The 80% confidence interval of prediction and average prediction results during 50 independent generations given the same input on ETTh1 dataset.:
@@ 6.3.2. Evaluation of Forecasting Uncertainty
In this section, we aim to evaluate TimeReasoner’s capability in forecasting uncertainty quantification, with the goal of examining how well it captures the inherent reasoning uncertainty of large language models during multi-step inference and assessing both the accuracy and reliability of its probabilistic outputs. As shown in Figure 5, predictions from the same LLM given identical input can vary notably, with the 80% prediction interval spanning a wide range across the forecast horizon. This highlights the inherent uncertainty in LLM-based forecasting and underscores the need to account for it when interpreting outputs. To quantify the reasoning uncertainty in slow-thinking forecasting, we compute the standard deviation of multiple predictions at each forecast step. Figure 6 shows per-step uncertainty curves across four datasets, each line representing the deviation trajectory over the forecast window. Uncertainty generally increases with forecast length, especially on ETTh1 and ETTh2, suggesting accumulation over longer generations. Sharp spikes—most notably on ETTh1 and ETTm1 reflect moments of high uncertainty. These patterns underscore the stochastic nature of LLM-based forecasting and the need for fine-grained uncertainty estimation.
@@ 6.4. Evaluation Under Different TSF Settings
In previous approaches, forecasting errors generally exhibit a positive correlation with the predict window length H since longer horizons introduce greater error propagation. Whether this principle still holds for TimeReasoner, which relies on step-by-step slow reasoning rather than learned representations, remains unclear. To investigate this, we fix the lookback window L = 96 and gradually increase the predict window $H\in\left\{48,72,96,120,144,168,192\right\}$. As shown in Figure 4, MSE increases with a longer predict window. This also proves that the previous principle is applicable to TimeReasoner. This performance degradation suggests that TimeReasoner’s reasoning becomes more difficult since it extrapolates further from observed data, limiting its effectiveness for long-term forecasting.
@@ 6.5. Hyperparameter Sensitivity Analysis
To address RQ5, we analyze the effects of inference-time hyperparameters, such as temperature and base LLM selection, on forecasting accuracy and stability.
@@
Temperature controls the randomness of token sampling during LLM generation, thus influencing the determinism and diversity of outputs. To examine its effect on TimeReasoner, we vary the decoding temperature and measure forecasting performance. As shown in Figure 7, the MSE follows a non-monotonic trend: starting from $T=0$, performance improves and reaches its best at $T=0.6$, after which it degrades. Low temperatures yield stable but rigid forecasts, often missing subtle temporal variations, while high temperatures inject excessive randomness, increasing variance and reducing reliability. The optimal point reflects a balance between stability and diversity in reasoning.
Figure 6. Standard deviation of predictions at each forecast step across ETTh1 and ETTh2 datasets.:
Figure 7. Sensitivity analysis of TimeReasoner to the temperature parameter across ETTh1, ETTh2, and ETTm2 datasets.:
Figure 8. Comparative performance of TimeReasoner with different base LLMs on Exchange and NASDAQ datasets.:
@@ 6.5.2. Performance Comparison across Base LLMs
We compare TimeReasoner with six base reasoning LLMs: DeepSeek-R1-0528, DeepSeek-R1, Gemini-2.5 Pro, QwQ-plus, DeepSeek-R1-Distill-Llama-70B, and GPT-5-Thinking. As shown in Figure 8, DeepSeek-R1-0528 achieves the highest accuracy on both Exchange and NASDAQ, followed closely by DeepSeek-R1, while the distilled variant performs worst, underscoring the need to retain model capacity and reasoning depth for temporal generalization. The improvement from DeepSeek-R1 to DeepSeek-R1-0528 further indicates that our framework can continue to benefit from future advances in base model reasoning. Interestingly, GPT-5-Thinking underperforms despite its size, often defaulting to code-style heuristics (e.g., linear or mean extrapolation) rather than attending to distributional cues in the lookback window. This causes bias accumulation, weaker responsiveness to regime shifts, and reduced accuracy. These results suggest that how an LLM reasons—remaining data-attentive versus shortcutting via tool-like computations—can matter more than parameter count, and that constraining such shortcuts with observation-first prompting and data-focused exemplars can improve robustness.
@@ 6.6. Visualization Analysis
To address RQ6, we analyze intermediate reasoning traces and a representative forecasting case to assess their coherence and alignment with the input data.
@@ 6.6.1. Reasoning Trace Analysis
Building on TimeReasoner’s strong zero-shot forecasting capabilities, we analyze its internal chain of thought (CoT), which qualitatively unfolds in three stages: (1) Data analysis and pattern recognition, where the model identifies structural features such as intervals, frequency, trends, and seasonality; (2) Forecasting strategy evaluation, where it considers methods ranging from seasonal baselines to trend extrapolation, reflecting on their trade-offs; and (3) Reflection and contextual assessment, where the model critiques its own reasoning, noting strengths like capturing seasonality and limitations such as lacking complex statistical mechanisms—demonstrating adaptive and context-aware thinking.
Figure 9. Heatmap of MSE rankings across CoT length groups on ETTh1 and ETTh2. The x-axis shows MSE ranks (#1–#10), and the y-axis shows CoT token deciles from shortest (bottom) to longest (top).:
@@ 6.6.2. Forecasting Case Analysis
Figure 10 presents an example of TimeReasoner’s inference-time reasoning, illustrating its slow-thinking capability for 96-96 time series forecasting. The model begins by identifying underlying temporal patterns such as trends and seasonality, followed by deliberation over suitable forecasting strategies. It then examines recent observations in detail, detecting potential anomalies and temporal shifts. Based on this analysis, TimeReasoner constructs a forecasting plan by aggregating patterns over comparable historical periods and ensuring continuity in the predicted sequence. The visualization on the right confirms that the predicted window preserves both seasonal and trend components from the lookback window, demonstrating the model’s ability to perform step-by-step reasoning akin to a human expert.
Figure 10. A case demonstrating how TimeReasoner performs slow-thinking capability for 96-96 time series reasoning.:
Figure 11. Four representative failure modes of TimeReasoner, including (a) Peak Clipping, (b) Phase-Shift Error, (c) Copy-Paste Repeat, and (d) Constant Collapse.:
@@ 6.6.3. Failure Case Analysis
These patterns reveal distinct weaknesses in amplitude capture, temporal alignment, adaptability, and output diversity. Such errors can severely undermine forecast reliability—misaligned peaks may cause decision failures in time-critical applications, excessive repetition can overlook regime changes or unexpected events, and constant predictions fail to reflect dynamic trends, leading to missed opportunities or inadequate responses. Beyond lowering pointwise accuracy, these issues also reduce the interpretability and trustworthiness of the model’s reasoning process. Addressing them will require enhancing sensitivity to extremes, improving phase anchoring to the correct temporal context, and discouraging trivial or degenerate outputs, thereby ensuring that the model remains both accurate and responsive across diverse forecasting scenarios.

### v4

@@ 5.3. Implementation Details
The TimeReasoner is model-independent; in this work, we adopt DeepSeek-R1222https://huggingface.co/deepseek-ai/DeepSeek-R1 for its open-source availability and ability to expose complete reasoning traces. By default, all experiments are conducted under the one-shot reasoning paradigm, with inference performed via the official API [^3] https://platform.deepseek.com using temperature=0.6, top-p=0.7, and max-tokens=8,192. Foundation model baselines follow the same zero-shot inference protocol, while deep learning baselines are trained using the Adam optimizer (kingma2014adam) in PyTorch (paszke2019pytorch) on a single NVIDIA GeForce RTX 4090D GPU, with official implementations and recommended parameters. The lookback window is set to $L=36$ for NASDAQ, $L=300$ for VitalDB, and $L=96$ for other datasets; the predict window is $H=36$, $H=200$, and $H=96$ respectively. We adopt a channel-independent evaluation. Following standard TSF practice, we report mean squared error (MSE) to measure squared deviations and mean absolute error (MAE) to measure absolute deviations between predictions and ground truth.
@@ 6.1. Forecasting Performance Comparison
To address RQ1, we evaluate TimeReasoner’s forecasting performance across multiple benchmarks against state-of-the-art baselines. Table 2 summarizes the forecasting results of TimeReasoner, demonstrating its competitive performance over other baseline models. Results on ten datasets show that TimeReasoner achieves competitive forecasting results on most datasets, especially on ETTh1 and AQShunyi. We observe that TimeReasoner particularly excels on datasets characterized by complex temporal dynamics. On datasets like AQWan, AQShunyi and Wind, TimeReasoner consistently surpasses other methods by effectively capturing underlying dependencies. These findings underscore TimeReasoner’s outstanding capability in tackling difficult forecasting tasks and its generalization to complex time series. We also observe that the model achieves performance comparable to strong baselines on other datasets. This suggests that the proposed approach does not introduce performance degradation outside scenarios involving complex temporal dynamics.
@@ 6.2. Analysis of Key Settings in TimeReasoner
To address RQ2, we conduct ablations on key components of TimeReasoner, including hybrid prompt constructions and reasoning strategies, to assess their impact on accuracy.
@@ 6.2.1. Impact of Hybrid Prompt
We study how prompt-design choices affect TimeReasoner’s forecasting performance by ablating three components: timestamps, contextual information, and normalization, as shown in Table 3. For timestamps, we consider two perturbations: removal, which discards all timestamp annotations, and shifting, which applies a uniform forward/backward offset while preserving relative intervals. Shifting results in only marginal degradation, suggesting robustness to changes in calendar-time anchoring, whereas removal causes a pronounced drop, highlighting the importance of temporal cues.
For contextual information, removing domain-specific descriptors degrades performance on datasets with strong seasonal or location patterns (e.g., Wind) but has little impact on highly volatile series (e.g., NASDAQ), indicating that the benefit of context is dataset-dependent. For normalization, we compare Z-score and RevIN with using raw time series and find that unnormalized inputs perform best. This suggests that standard normalization, while often helpful for deep forecasting models, can attenuate salient structures (e.g., trends and abrupt changes) that are useful for slow-thinking reasoning, and that preserving the original scale may be more compatible with our prompting setup.
Figure 2. Comparison of TimeReasoner’s forecasting capabilities under the three reasoning strategies with experiments conducted on ETTh2 (left) and Exchange (right) datasets.
Comparison of TimeReasoner’s forecasting capabilities under the three reasoning strategies with experiments conducted on ETTh2 (left) and Exchange (right) datasets.
@@ 6.2.2. Impact of Various Reasoning Strategies
We hypothesize that Rollout Reasoning may be more effective for short-horizon forecasting, as multi-round prediction can limit error propagation when the horizon is small, while Decoupled Reasoning may be better suited for long-horizon forecasting by enabling more thorough refinement within a single inference pass. We fix the lookback window to $L=96$ and evaluate four predict windows $H\in\{48,96,144,192\}$.
Figure 2 shows that One-Shot Reasoning remains competitive across all horizons. Rollout Reasoning performs best at short horizons ($H=48,96$), but its error grows more rapidly as the horizon increases ($H=144,192$), consistent with error accumulation from iterative rollouts. In contrast, Decoupled Reasoning becomes the best-performing method at longer horizons, suggesting that its within-pass refinement yields more stable predictions when long-range dependencies dominate.
Table 4. Experimental results on ETTh1 and ETTm1 datasets showing model stability under different input settings.
@@
Figure 3. Relationship between TimeReasoner’s forecasting performance and lookback window and predict window.
Relationship between TimeReasoner’s forecasting performance and lookback window and predict window.
@@ 6.3. Trustworthiness Evaluation
To address RQ3, we assess TimeReasoner’s robustness by simulating missing and noisy inputs, and further evaluate its uncertainty estimation capability.
6.3.1. Stability Evaluation w.r.t. Missing Input
We evaluate the stability of TimeReasoner under missing values, a common challenge in real-world time series (e.g., due to sensor failures (che2018recurrent)). While traditional models typically rely on imputation, slow-thinking LLM-based approaches can process raw, incomplete data directly. We consider three variants: (1) No Imputation (No-Imp), where missing entries are removed; (2) None Placeholder Imputation (None-Imp), where missing values are replaced with a None token; and (3) Linear Interpolation Imputation (Lin-Imp) (chow1971best). As shown in Table 4, TimeReasoner performs best with fully observed data, demonstrating its ability to capture complete temporal dependencies. It remains robust under linear interpolation, while using None placeholders causes moderate degradation due to increased uncertainty. In contrast, dropping missing entries leads to a sharp decline, as it disrupts the temporal structure. These results suggest that preserving temporal continuity is more beneficial than deleting missing entries even when using placeholders.
Figure 4. The 80% confidence interval of prediction and average prediction results during 50 independent generations given the same input on ETTh1 dataset.
The 80% confidence interval of prediction and average prediction results during 50 independent generations given the same input on ETTh1 dataset.
@@ 6.3.2. Evaluation of Forecasting Uncertainty
In this section, we aim to evaluate TimeReasoner’s capability in forecasting uncertainty quantification, with the goal of examining how well it captures the inherent reasoning uncertainty of large language models during multi-step inference and assessing both the accuracy and reliability of its probabilistic outputs.
As shown in Figure 4, predictions from the same LLM given identical input can vary notably, with the 80% prediction interval spanning a wide range across the forecast horizon. This highlights the inherent uncertainty in LLM-based forecasting and underscores the need to account for it when interpreting outputs. To quantify the reasoning uncertainty in slow-thinking forecasting, we compute the standard deviation of multiple predictions at each forecast step. Figure 5 shows per-step uncertainty curves across four datasets, each line representing the deviation trajectory over the forecast window. Uncertainty generally increases with forecast length, especially on ETTh1 and ETTh2, suggesting accumulation over longer generations. Sharp spikes—most notably on ETTh1 and ETTm1 reflect moments of high uncertainty. These patterns underscore the stochastic nature of LLM-based forecasting and the need for fine-grained uncertainty estimation.
@@ 6.4. Evaluation Under Different TSF Settings
In previous approaches, forecasting errors generally exhibit a positive correlation with the predict window length H since longer horizons introduce greater error propagation. Whether this principle still holds for TimeReasoner, which relies on step-by-step slow reasoning rather than learned representations, remains unclear. To investigate this, we fix the lookback window L = 96 and gradually increase the predict window $H\in\left\{48,72,96,120,144,168,192\right\}$. As shown in Figure 3, MSE increases with a longer predict window. This suggests that the previous principle is applicable to TimeReasoner. This performance degradation suggests that TimeReasoner’s reasoning becomes more difficult since it extrapolates further from observed data, limiting its effectiveness for long-term forecasting.
Figure 5. Standard deviation of predictions at each forecast step across ETTh1 and ETTh2 datasets.
Standard deviation of predictions at each forecast step across ETTh1 and ETTh2 datasets.
@@ 6.5. Hyperparameter Sensitivity Analysis
To address RQ5, we analyze the effects of inference-time hyperparameters, such as temperature and base LLM selection, on forecasting performance and stability.
@@
Temperature controls the randomness of token sampling during LLM decoding, thereby affecting output determinism and variability. To quantify its impact on TimeReasoner, we vary the decoding temperature while keeping the prompt construction and all other settings (including the lookback window and predict window) fixed, and evaluate forecasting performance. As shown in Figure 6, MSE exhibits a non-monotonic pattern: starting from $\tau=0$, performance improves and achieves the lowest MSE at $\tau=0.6$, but degrades as $\tau$ increases further. Lower temperatures yield more stable yet less flexible generations that may miss subtle temporal variations, whereas higher temperatures introduce excessive randomness, increasing variance and reducing reliability. These results suggest that an intermediate temperature provides a better trade-off between stability and controlled diversity in generation.
Figure 6. Sensitivity analysis of TimeReasoner to the temperature parameter across ETTh1, ETTh2, and ETTm2 datasets.
Sensitivity analysis of TimeReasoner to the temperature parameter across ETTh1, ETTh2, and ETTm2 datasets.
Figure 7. Comparative performance of TimeReasoner with different base LLMs on Exchange and NASDAQ datasets.
Comparative performance of TimeReasoner with different base LLMs on Exchange and NASDAQ datasets.
@@ 6.5.2. Performance Comparison across Base LLMs
We compare TimeReasoner across six backbone reasoning LLMs: DeepSeek-R1-0528, DeepSeek-R1, Gemini-2.5 Pro, QwQ-plus, DeepSeek-R1-Distill-Llama-70B, and GPT-5-Thinking. As shown in Figure 7, DeepSeek-R1-0528 delivers the best forecasting performance on both Exchange and NASDAQ, followed closely by DeepSeek-R1, while the distilled variant performs worst, highlighting the performance gap introduced by distillation. The improvement from DeepSeek-R1 to DeepSeek-R1-0528 further indicates that TimeReasoner benefits from stronger backbone reasoning capabilities.
Interestingly, GPT-5-Thinking underperforms in this setting. Qualitative inspection suggests that it sometimes relies on simple code-style extrapolation heuristics (e.g., linear or mean extrapolation) instead of closely tracking distributional cues in the lookback window, which can reduce responsiveness to regime shifts and degrade accuracy. Overall, these results suggest that the style of reasoning—data-attentive inference versus heuristic shortcutting—can be as important as model scale, and that observation-first prompting and data-focused exemplars may be a promising direction for improving robustness.
@@ 6.6. Visualization Analysis
To address RQ6, we analyze intermediate reasoning traces and a representative forecasting case to assess whether the output is aligned with the input data.
@@ 6.6.1. Reasoning Trace Analysis
Building on TimeReasoner’s zero-shot forecasting performance, we examine its model-generated CoT-style reasoning traces. We observe a common three-stage structure: (1) Data inspection and pattern identification, where the model summarizes salient temporal characteristics (e.g., periodicity/seasonality, trend, and level shifts); (2) Strategy selection and justification, where it evaluates candidate forecasting heuristics (from seasonal baselines to trend extrapolation) and discusses their trade-offs; and (3) Self-reflection and contextual checks, where it critiques the chosen strategy, noting strengths (e.g., capturing seasonality) and limitations (e.g., lacking explicit statistical mechanisms). Overall, these traces suggest a diagnostic–select–verify workflow prior to producing the final forecast.
Figure 8. Heatmap of MSE rankings across CoT length groups on ETTh1 and ETTh2. The x-axis shows MSE ranks (#1–#10), and the y-axis shows CoT token deciles from shortest (bottom) to longest (top).
Heatmap of MSE rankings across CoT length groups on ETTh1 and ETTh2. The x-axis shows MSE ranks (#1–#10), and the y-axis shows CoT token deciles from shortest (bottom) to longest (top).
@@ 6.6.2. Forecasting Case Analysis
Figure 9 presents a representative example of TimeReasoner’s inference-time reasoning, illustrating its deliberative inference behavior for time-series forecasting with lookback window $L=96$ and predict window $H=96$. The model first summarizes salient temporal patterns (e.g., trend and seasonality) and then evaluates candidate forecasting strategies. It subsequently inspects the most recent observations, noting potential anomalies and regime shifts. Based on these cues, TimeReasoner forms a forecasting plan by referencing comparable historical periods and enforcing continuity across the predicted sequence. The visualization on the right shows that the predicted window preserves both seasonal and trend components observed in the lookback window, suggesting that the model can produce interpretable, step-by-step forecasts.
Figure 9. A case demonstrating how TimeReasoner performs slow-thinking capability for 96-96 time series reasoning.
A case demonstrating how TimeReasoner performs slow-thinking capability for 96-96 time series reasoning.
Figure 10. Four representative failure modes of TimeReasoner, including (a) Peak Clipping, (b) Phase-Shift Error, (c) Copy-Paste Repeat, and (d) Constant Collapse.
Four representative failure modes of TimeReasoner, including (a) Peak Clipping, (b) Phase-Shift Error, (c) Copy-Paste Repeat, and (d) Constant Collapse.
@@ 6.6.3. Failure Case Analysis
These patterns reveal distinct weaknesses in amplitude capture, temporal alignment, adaptability, and output diversity. Such errors can severely undermine forecast reliability—misaligned peaks may cause decision failures in time-critical applications, excessive repetition can overlook regime changes or unexpected events, and constant predictions fail to reflect dynamic trends, leading to missed opportunities or inadequate responses. Beyond lowering pointwise accuracy, these issues also reduce the explainability and trustworthiness of the model’s reasoning process. Addressing them will require enhancing sensitivity to extremes, improving phase anchoring to the correct temporal context, and discouraging trivial or degenerate outputs, thereby ensuring that the model remains both accurate and responsive across diverse forecasting scenarios.